OpenStudy (anonymous):

The x-intercept of the line whose equation is 2x - 3y = 6 is 3 6 -2

OpenStudy (anonymous):

When the line intercepts the x-axis, y must equal zero. So set y = 0 and solve for x. 2x - 3y = 6 2x - 3(0) = 6 2x = 6 x = 6/2 x = 3